OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the diagonal of a square whose vertices are A(-3, 3), B(3, 3), C(3, -3), and D(-3, -3). Find two equations, one for each diagonal. please help!! some one !!

OpenStudy (anonymous):

help

OpenStudy (shamim):

i can solve all of ur this kind of math. but now i hv to go out

OpenStudy (shamim):

i feel very good by solving ur math

OpenStudy (anonymous):

can you solve them ?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

some one help im still stuck xO

OpenStudy (anonymous):

i cant see what you attached

OpenStudy (anonymous):

Click on the attached file to download it and then check it out.

OpenStudy (shamim):

the equation of diagonal AC is\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{1})\]

OpenStudy (shamim):

here\[x _{1}=-3,y _{1}=3,x _{2}=3,y _{2}=-3\]

OpenStudy (shamim):

\[y-3=\frac{ -3-3 }{ 3-(-3) }(x-(-3))\]

OpenStudy (shamim):

\[y-3=\frac{ -6 }{ 6 }(x+3)\]

OpenStudy (shamim):

\[y-3=-x-3\]

OpenStudy (shamim):

\[x+y=0\]

OpenStudy (shamim):

now u can try to find out the equation of another diagonal BD